# Solving the Bongard-logo Problem by Modeling a Probabilistic Model

### v3

@@ Abstract
Abstract reasoning problems challenge the perceptual and cognitive abilities of AI algorithms, demanding deeper pattern discernment and inductive reasoning beyond explicit image features. This study introduces PMoC, a tailored probability model for the Bongard-Logo problem, achieving high reasoning accuracy by constructing independent probability models. Additionally, we present Pose-Transformer, an enhanced Transformer-Encoder designed for complex abstract reasoning tasks, including Bongard-Logo, RAVEN, I-RAVEN, and PGM. Pose-Transformer incorporates positional information learning, inspired by capsule networks’ pose matrices, enhancing its focus on local positional relationships in image data processing. When integrated with PMoC, it further improves reasoning accuracy. Our approach effectively addresses reasoning difficulties associated with abstract entities’ positional changes, outperforming previous models on the OIG, D3×3 subsets of RAVEN, and PGM databases. This research contributes to advancing AI’s capabilities in abstract reasoning and cognitive pattern recognition.
@@
Abstract reasoning, RPM problem, Bongard-logo problem.
@@ II-D Capsule Network
II-E Transformer
The Transformer[4] architecture is a distinctive deep learning framework that revolves around the self-attention mechanism to capture global dependencies within input sequences. Comprising of an encoder and a decoder, both structured with stacked layers of multi-head self-attention and feed-forward neural networks, the Transformer offers a paradigm shift in sequential data processing.
Initially, the input sequence undergoes an embedding layer transformation into fixed-dimensional vectors, augmented with positional encoding to retain positional information within the sequence. Subsequently, the encoder, composed of multiple identical layers, each encompassing two sub-layers, commences its operation. The first sub-layer, the multi-head self-attention layer, calculates attention scores between different positions in the input sequence, capturing global dependencies. The second sub-layer comprises a simple fully connected feed-forward neural network. Both sub-layers are followed by layer normalization operations and residual connections, ensuring model stability and training efficiency.
The decoder section, also comprising multiple identical layers, introduces an additional encoder-decoder multi-head attention layer in each decoding layer. This attention layer enables the decoder to attend to the encoder’s output during the generation of the output sequence, facilitating interaction between the encoder and decoder. Furthermore, the self-attention layers in the decoder incorporate a masking mechanism to prevent consideration of future information when generating outputs for the current position.
The uniqueness of the Transformer architecture lies in its exclusive reliance on attention mechanisms, circumventing the inherent sequential computation limitations of traditional recurrent and convolutional neural networks. This design allows for parallel processing of input sequences, significantly enhancing training speed and model performance. Additionally, the Transformer exhibits impressive scalability, accommodating tasks of varying scales and complexities through the addition of encoder and decoder layers and adjustment of model parameters.
In summary, the Transformer architecture’s innovative approach to sequence modeling, characterized by its reliance on self-attention mechanisms and parallel processing capabilities, has ushered in a new era of advancements in natural language processing and beyond.
II-F Vision Transformer
Vision Transformer (ViT)[34] is a groundbreaking image classification model that adapts the Transformer architecture from the domain of Natural Language Processing (NLP) and applies it successfully to Computer Vision (CV) tasks. The uniqueness of ViT lies in its treatment of images as sequential data, akin to text sequences in NLP, thereby leveraging the powerful sequence modeling capabilities of the Transformer.
In ViT, the input image is initially divided into fixed-size patches, which are then flattened and transformed into feature vectors through linear transformations. These feature vectors are augmented with positional encodings to retain their spatial positional information within the image. Subsequently, the processed feature vectors are fed into a standard Transformer encoder for processing.
The Transformer encoder comprises multiple layers of self-attention and feed-forward neural networks, arranged in a stacked manner. The self-attention layers capture dependencies between different positions within the input sequence, while the feed-forward neural networks introduce non-linear transformations to enhance the model’s expressive power. Residual connections and layer normalization techniques are employed in each layer of the encoder to ensure model stability and convergence speed.
Ultimately, the output from the Transformer encoder is fed into a fully connected layer for classification, yielding the predicted categories for the image. Due to its sequence-to-sequence modeling approach similar to NLP, ViT effectively handles images of various sizes and resolutions and exhibits impressive scalability.
In summary, by adapting the Transformer architecture from NLP, Vision Transformer (ViT) has ushered in a new era of advancements in Computer Vision. Its emergence has not only challenged the dominance of traditional Convolutional Neural Networks (CNNs) in image classification tasks but has also laid a solid foundation for the development of subsequent visual Transformer models.
II-G Sinkhron distance
@@
III Methodology
The Bongard-logo[16] and Raven’s Progressive Matrices (RPM)[15] tasks are both paradigms of abstract reasoning that necessitate the extractor’s adeptness in unveiling precise concepts veiled within abstract representations. These concepts are indicative of more elevated abstractions, surpassing rudimentary ideas such as “pixel configuration patterns”. Frequently, these elevated abstractions encapsulate diverse degrees of human-centered presupposed knowledge, including elements like form, dimension, hue, spatial orientations, concave/convex contours of shapes, and shape completion. Through the employment of distinct querying approaches, the Bongard-logo and RPM tasks endeavor to evaluate the extractor’s expertise in unveiling and assimilating these abstract concepts.
@@ III-A A method for solving the bongard-logo problem based on sinkhorn distance (SBSD)
Therefore, we aim to devise a deep learning algorithm, denoted as $f_{\theta}(z|x)$, which can map samples $x_{ij}$ to latent variables $z_{ij}$. We hope that $f_{\theta}(z|x)$ can serve as a distribution transformation function, and by measuring the distribution of the latent representation, we can address the Bongard-Logo problem. Thus, the key objective is to ensure that the cross-measure between the latent variable distribution of the primary group, denoted as $p_{i}^{\prime}(z|y)$, and the latent variable distribution of the auxiliary group, denoted as $q_{i}^{\prime}(z|y)$, is significantly low. The Kullback-Leibler (KL) divergence, a measure used to quantify the similarity between distributions, can be employed for this purpose[7]. The formula for calculating KL divergence is provided below.
@@
Due to the characteristics of small sample learning problems, it is challenging to estimate the distributions $p_{i}^{\prime}(z|y)$ and $q_{i}^{\prime}(z|y)$ of Bongard-Logo image representations. Additionally, the limitations of the Kullback-Leibler (KL) divergence make it difficult to directly optimize for these conditions and obtain a highly effective deep model.
Hence, we endeavored to employ baseline models, such as ResNet18, for encoding images within the Bongard-Logo paradigm $\{x_{ij}\}_{j=1}^{14}$ into meaningful representations $\{z_{ij}\}_{j=1}^{14}$. These representations, within a Bongard-Logo question, were treated as samples drawn from $p_{i}^{\prime}(z|y)$ and $q_{i}^{\prime}(z|y)$ distributions. Subsequently, we imposed constraints on the Sinkhorn distance[35] between $p_{i}^{\prime}(z|y)$ and $q_{i}^{\prime}(z|y)$ through these representations. This approach was chosen because the Sinkhorn algorithm calculates the distance without requiring knowledge of the specific forms of the distributions, thus making it ideally suited for scenarios where the forms of $p_{i}^{\prime}(z|y)$ and $q_{i}^{\prime}(z|y)$ are unknown. In the detail, we randomly divided the representations within the primary group $\{z_{ij}\}_{j=1}^{7}$ into two distinct subgroups: $\{z_{it}\}$ and $\{{z}_{i\tilde{t}}\}$. By leveraging the Sinkhorn distance between these subgroups, we imposed constraints to enforce the representations in the primary group $\{z_{ij}\}_{j=1}^{7}$ to follow a consistent, albeit unknown, distribution. Similarly, we introduced analogous constraints between the distributions denoted as $p_{i}^{\prime}(z|y)$ and $q_{i}^{\prime}(z|y)$, utilizing representations from both the primary and auxiliary groups to minimize the cross-measure. The aforementioned constraints can be concretely expressed as mathematical loss function:
@@
In summary, SBSD treats the representations of images within the primary group as samples from the primary group distribution $p_{i}^{\prime}(z|y)$, and the auxiliary group follows the same approach. By utilizing these samples, the Sinkhorn distance is employed to constrain the distance between the primary $p_{i}^{\prime}(z|y)$ and auxiliary distributions $q_{i}^{\prime}(z|y)$.
@@ III-B PMoC(Probability Model of Concept)
Our PMoC consists of two components: module $f_{\theta}(z|x)$ and module $g_{\omega}(\mu,\sigma^{2}|z)$. The module $f_{\theta}(z|x)$ maps the input $x$ to its latent representation $z$, while $g_{\omega}(\mu,\sigma^{2}|z)$ is utilized to compute the $\mu$ and $\sigma^{2}$ for Gaussian distribution $p^{\prime}_{i}(z|y)$.
@@
We present an approach that combines ResNet18 with Transformer-Encoder, employing this integration as the architecture for the function $f_{\theta}(z_{ij}|x_{ij})$. In detail, we encode the Bongard-Logo problem images $x_{ij}$ into feature maps $h_{ij}\in R^{h\times w\times d}$ using ResNet18. Subsequently, we calculatethe self-attention results $z^{\prime}_{ij}\in R^{n\times d}$ among all receptive fields within the feature maps $h_{ij}$. This computation enables the $n$ attention outputs $z^{\prime}_{ij}$ to capture global information of the image in varying degrees, ensuring that each output retains some level of contextual understanding. We encoded Bongard-Logo images from multiple perspectives using the aforementioned method. We denote each individual perspective within $z^{\prime}_{ij}$ as $z_{ij}\in R^{d}$. The feedforward process of the network $f_{\theta}(z_{ij}|x_{ij})$ is illustrated in the figure 5.
@@
The network $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ is much more of its name. It encodes the mean and variance of the $i$-th primary group distribution based on the latent representations of the $i$-th primary group samples, and $i$ still represents the problem index in Bongard-Logo. Specifically, we uniformly process each perspective $z_{ij}$ within $z^{\prime}_{ij}$. Under a single perspective, we treat $\{z_{ij}\}^{6}_{j=1}$ as six tokens and, after appending two optimizable vectors along with positional embedding, input them into $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ with a Transformer-Encoder backbone for computing attention results. We extract two optimizable vectors from the attention results, serving as the mean $\mu_{i}\in R^{d}$ and logarithm of variance $\log(\sigma_{i}^{2})\in R^{d}$ for a multivariate Gaussian distribution. The encoding distribution pattern of module $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ follows a similar approach to that of VAE[7]; however, it does not incorporate variational inference in its design. The aforementioned process parameterizes $p_{i}^{\prime}(z|y)$ as a multivariate Gaussian distribution and outlines the computation of $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$. We feed 8 vectors $\{z_{ij^{\prime}}\}^{14}_{j^{\prime}=7}$ into this distribution for calculating 8 logical probabilities $\{{p}_{i}^{\prime}(z_{ij^{\prime}}|\{z_{ij}\}^{6}_{j=1},\omega)\}_{j^{\prime}=7}^{14}$ in one perspective. In this paper, we calculate the final probability ${p}_{i}^{\prime}(z^{\prime}_{ij^{\prime}})$ by averaging the logical probabilities ${p}_{i}^{\prime}(z_{ij^{\prime}}|\{z_{ij}\}^{6}_{j=1},\omega)$ obtained from all perspectives. By utilizing the cross entropy loss function, we constrain the final probability ${p}_{i}^{\prime}(z^{\prime}_{i7})$ to approach 1 while ensuring other probabilities $\{{p}_{i}^{\prime}(z^{\prime}_{ij})\}_{j=8}^{14}$ are close to 0. Figure 6 illustrates in detail the forward process of the network $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ under a single perspective in $z_{ij}\in R^{n\times d}$. However, this approach did not yield satisfactory results.
@@
We speculate that the reason for the ineffectiveness of the aforementioned method lies in the unknown true form of the distribution $p_{i}^{\prime}(z|y)$, which may be a complex mixed Gaussian distribution. Therefore, it may not be reasonable to parameterize $p_{i}^{\prime}(z|y)$ as a multivariate Gaussian distribution and compute the probability $p_{i}^{\prime}(z_{ij^{\prime}}|y)$. Additionally, the formula for calculating the logical probability of a vector $z$ belonging to the parameterized distribution is provided as follows.
@@
Where $m$ represents the dimension index of the vector. $\mu$ represents the mean of the multivariate Gaussian distribution, $\sigma$ represents the standard deviation of the multivariate Gaussian distribution, and $z$ is the representation of some unknown data. It is evident that optimizing the network using the given formula (3) as the loss function presents challenges, as optimizing the above equation can be indirectly viewed as optimizing the Mean Squared Error (MSE) loss between vectors. Moreover, these vectors have not undergone normalization or squashing, making it difficult to optimize the network from the perspective of implementation. Consequently, in this paper, we train a deep network to directly fit probability $p_{i}^{\prime}(z_{ij^{\prime}}|y)$. Inevitably the network’s formulation has been transformed from $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ to $g_{\omega}(p_{i}^{\prime}(z_{ij^{\prime}}|y)|\{z_{ij}\}^{6}_{j=1},z_{ij^{\prime}})$, thereby simplifying the problem.
Different from previous approaches $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ that fit the distribution $p_{i}^{\prime}(z|y)$, this paper employs a Transformer-Encoder to fit the probability $p_{i}^{\prime}(z_{ij^{\prime}}|y)$ directly, which enables the direct computation of the probability that $z_{ij^{\prime}}$ belongs to the distribution $p_{i}^{\prime}(z|y)$.
@@
Specifically, in this paper, we maintain an egalitarian approach towards each perspective $z_{ij}\in R^{d}$ within the outcomes $z^{\prime}_{ij}\in R^{n\times d}$ of $f_{\theta}(z_{ij}|x_{ij})$. Within the confines of a singular perspective, we amalgamate $\{z_{ij}\}^{6}_{j=1}$ with one of the pending representations $\{z_{ij^{\prime}}\}^{14}_{j^{\prime}=7}$, resulting in a set of seven tokens. Following the embedding of positional encoding, these tokens are then systematically fed into a Transformer-Encoder, aimed at yielding 7 attention results. All 7 attention results are mapped to 7 probabilities by a MLP. The mean of the mapped outcomes is considered as the logical probability ${p}_{i}^{\prime}(z_{ij^{\prime}}|y)$ corresponding to the specified perspective. This constitutes the feedforward process of the new $g_{\omega}(p_{i}^{\prime}(z_{ij^{\prime}}|y)|\{z_{ij}\}^{6}_{j=1},z_{ij^{\prime}})$. Each pending representation $z_{ij^{\prime}}$, receives a corresponding probability output $p_{i}^{\prime}(z_{ij^{\prime}}|y)$. Each logical probabilities $\{{p}_{i}^{\prime}(z_{ij^{\prime}}|y)\}_{j^{\prime}=7}^{14}$ are averaged to yield the final probability value $\{p_{i}^{\prime}(z^{\prime}_{ij^{\prime}}|y)\}_{j^{\prime}=7}^{14}$. In this paper, we still optimize the model using the cross-entropy loss function. Figure 7 visually depicts the process of $g_{\omega}(p_{i}^{\prime}(z_{ij^{\prime}}|y)|\{z_{ij}\}^{6}_{j=1},z_{ij^{\prime}})$ handling a single perspective.
@@
In summary, this paper makes two compromises for PMoC. Firstly, we acknowledge that compared to $p_{i}^{\prime}(z)$, $q_{i}^{\prime}(z)$ is more likely to be a complex distribution. For instance, $p_{i}^{\prime}(z)$ may follow a Gaussian distribution, while $q_{i}^{\prime}(z)$ may adhere to a mixed Gaussian distribution. Consequently, we shift our focus to estimating the distribution $p_{i}^{\prime}(z)$ and utilize the probability of a representation belonging to this distribution to determine whether it is a primary or auxiliary group representation. Secondly, due to the challenges associated with optimizing the model from the perspective of estimating $p_{i}^{\prime}(z)$, this paper opts to directly compute the probability of a representation belonging to the distribution $p_{i}^{\prime}(z)$.
IV Pose-tansformer

### v7

@@ Abstract
Abstract reasoning problems pose challenges to the perception and cognition abilities of AI algorithms, demanding deeper pattern recognition and inductive reasoning beyond mere identification of explicit image features. In this study, we introduce PMoC, a probabilistic model tailored for the Bongard-Logo problem, achieving high reasoning accuracy through the construction of an conditional probabilistic model. Additionally, we have designed the Pose-Transformer, an enhanced Transformer-Encoder specifically crafted for complex abstract reasoning tasks, including Bongard-Logo, RAVEN, I-RAVEN, and PGM. Inspired by the pose matrix in capsule networks, Pose-Transformer strengthens the focus on positional relationships between local features when processing image data. When combined with PMoC, it can further enhance reasoning accuracy. Our Pose-Transformer effectively addresses reasoning difficulties associated with changes in the position of abstract entities, outperforming previous models on RAVEN’s OIG, D3x3 subsets, and the PGM dataset. Finally, considering the deployment difficulties arising from the large number of Pose-Transformer parameters, this paper presents a lightweight version, Straw-Pose-Transformer, which maintains performance while significantly reducing the parameter count. This study contributes to enhancing AI capabilities in abstract reasoning and cognitive pattern recognition.
@@
Abstract reasoning, Raven’s Progressive Matrices, Bongard-logo problem, Probabilistic Model.
@@ II-D Capsule Network
II-E Transformer and Vision Transformer (ViT)
The Transformer [4] model, originally designed for machine translation tasks, is a encoder-decoder architecture that relies heavily on attention mechanisms, particularly self-attention. This mechanism enables the model to simultaneously consider other words in a sentence while processing a specific word, effectively capturing contextual relationships between them. On the other hand, Vision Transformer (ViT) [34] represents the application of the Transformer model in the image domain. It processes image data by dividing it into small patches, linearly embedding each patch into a vector, and then leveraging attention mechanisms to capture global contextual information for image recognition and processing. Both Transformer and Vision Transformer utilize self-attention mechanisms and encoder-decoder architectures, but they differ in the types of data they handle and the manner in which they process it.
II-F Sinkhron distance
@@
III Methodology for Bongard-Logo
The Bongard-logo[16] and Raven’s Progressive Matrices (RPM)[15] tasks are both paradigms of abstract reasoning that necessitate the extractor’s adeptness in unveiling precise concepts veiled within abstract representations. These concepts are indicative of more elevated abstractions, surpassing rudimentary ideas such as “pixel configuration patterns”. Frequently, these elevated abstractions encapsulate diverse degrees of human-centered presupposed knowledge, including elements like form, dimension, hue, spatial orientations, concave and convex contours of shapes, and shape completion. Through the employment of distinct querying approaches, the Bongard-logo and RPM tasks endeavor to evaluate the extractor’s expertise in unveiling and assimilating these abstract concepts.
This paper attempts to propose a methodology that relies on analyzing the distribution of images in the primary and auxiliary groups of Bongard-Logo problems to solve them. This methodology aims to provide a new direction different from the previous ones for solving Bongard-Logo problems.
@@ III-A A method for solving the bongard-logo problem based on sinkhorn distance (SBSD)
Directly deriving the image distributions $p_{i}(x|y)$ and $q_{i}(x|y)$ of Bongard-Logo without dimensionality reduction and computing the cross measure between these two distributions is quite challenging. Moreover, such an approach would also result in a waste of resources and time. Therefore, we aim to devise a deep learning algorithm, denoted as $f_{\theta}(z|x)$, which can map samples $x_{ij}$ to latent variables $z_{ij}$. We hope that $f_{\theta}(z|x)$ can serve as a distribution transformation function, and by measuring the distribution of the latent representation, we can address the Bongard-Logo problem. Accordingly, the key objective is to ensure that the encoder $f_{\theta}(z|x)$ possesses a specific image encoding capability, which entails producing encoded latent variables that demonstrate specific characteristics. Specifically, the distribution of latent variables within the primary group (represented as $p_{i}^{\prime}(z|y)$) and the distribution of latent variables within the auxiliary group (represented as $q_{i}^{\prime}(z|y)$) should exhibit very low cross-measure. The Kullback-Leibler (KL) divergence, a measure used to quantify the similarity between distributions, can be employed for this purpose[7]. The formula for calculating KL divergence is provided below.
@@
However, due to the characteristics of small sample learning problems, it is challenging to estimate the distributions $p_{i}^{\prime}(z|y)$ and $q_{i}^{\prime}(z|y)$ of Bongard-Logo image representations. Estimating the specific forms of distributions $p_{i}^{\prime}(z|y)$ and $q_{i}^{\prime}(z|y)$ poses an even greater challenge. Additionally, the limitations of the Kullback-Leibler (KL) divergence make it difficult to directly optimize for these conditions and obtain a highly effective deep model.
Hence, we endeavored to utilize baseline models, specifically ResNet18, for the purpose of encoding images within the Bongard-Logo paradigm, denoted as $\{x_{ij}\}_{j=1}^{14}$, into meaningful representations, designated as $\{z_{ij}\}_{j=1}^{14}$. These representations, in the context of a Bongard-Logo question, were interpreted as samples drawn from the distributions $p_{i}^{\prime}(z|y)$ and $q_{i}^{\prime}(z|y)$. Subsequently, constraints were imposed on the Sinkhorn distance between $p_{i}^{\prime}(z|y)$ and $q_{i}^{\prime}(z|y)$ through these representations. This methodology was designed primarily because the Sinkhorn algorithm computes the distance without requiring explicit knowledge of the distributions’ specific forms, making it highly suitable for scenarios where the forms of $p_{i}^{\prime}(z|y)$ and $q_{i}^{\prime}(z|y)$ remain unknown.
Specifically, the representations within the primary group, denoted as $\{z_{ij}\}_{j=1}^{7}$, were randomly partitioned into two distinct subgroups: $\{z_{it}\}$ and $\{{z}_{i\tilde{t}}\}$. By utilizing the Sinkhorn distance between these subgroups, constraints were imposed to ensure that the representations within the primary group $\{z_{ij}\}_{j=1}^{7}$ adhered to a consistent, albeit unknown, distribution. Analogously, similar constraints were introduced between the distributions represented by $p_{i}^{\prime}(z|y)$ and $q_{i}^{\prime}(z|y)$, leveraging representations from both the primary and auxiliary groups to minimize the cross-measure.
The aforementioned constraints can be explicitly articulated through a mathematical loss function.
@@
In summary, SBSD treats the representations of images within the primary group as samples from the primary group distribution $p_{i}^{\prime}(z|y)$, and the auxiliary group follows the same approach. By utilizing these samples, the Sinkhorn distance is employed to calculate and represent the distance between the primary $p_{i}^{\prime}(z|y)$ and auxiliary distributions $q_{i}^{\prime}(z|y)$. This distance can be used to solve the Bongard-Logo problem.
@@ III-B PMoC(Probability Model of Concept)
Our PMoC consists of two components: module $f_{\theta}(z|x)$ and module $g_{\omega}(\mu,\sigma^{2}|z)$. The module $f_{\theta}(z|x)$ maps the input $x$ to its latent representation $z$, while $g_{\omega}(\mu,\sigma^{2}|z)$ is utilized to compute the $\mu$ and $\sigma^{2}$ for the Gaussian distribution $p^{\prime}_{i}(z|y)$.
@@
Regarding the network implementation of $f_{\theta}(z_{ij}|x_{ij})$, we present an approach that combines ResNet18 with Transformer-Encoder, employing this integration as the architecture for the function $f_{\theta}(z_{ij}|x_{ij})$. In detail, we encode the Bongard-Logo problem images $x_{ij}$ into feature maps $h_{ij}\in R^{h\times w\times d}$ using ResNet18. Subsequently, we calculatethe self-attention results $z^{\prime}_{ij}\in R^{n\times d}$ among all receptive fields within the feature maps $h_{ij}$. This computation enables the $n$ attention outputs $z^{\prime}_{ij}$ to capture global information of the image in varying degrees, ensuring that each output retains some level of contextual understanding. We encoded Bongard-Logo images from multiple scales and perspectives using the aforementioned method. We denote each individual perspective within $z^{\prime}_{ij}$ as $z_{ij}\in R^{d}$. Since $g_{\omega}(\cdot|\cdot)$ in this paper treats each perspective within $z^{\prime}_{ij}$ equally and in parallel, we do not assign different notations to individual perspectives within $z^{\prime}_{ij}$ and uniformly denote them as $z_{ij}$. The feedforward process of the network $f_{\theta}(z_{ij}|x_{ij})$ is illustrated in the figure 5.
@@
The network $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ is much more of its name. It encodes the mean and variance of the $i$-th primary group distribution based on the latent representations of the $i$-th primary group samples, and $i$ still represents the problem index in Bongard-Logo. Specifically, we uniformly process each perspective $z_{ij}$ within $z^{\prime}_{ij}$. Under a single perspective, we treat $\{z_{ij}\}^{6}_{j=1}$ as six tokens and, after appending two optimizable vectors along with positional embedding, input them into $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ with a Transformer-Encoder backbone for computing attention results. We extract two optimizable vectors from the attention results, serving as the mean $\mu_{i}\in R^{d}$ and logarithm of variance $\log(\sigma_{i}^{2})\in R^{d}$ for a multivariate Gaussian distribution. The encoding distribution pattern of module $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ follows a similar approach to that of VAE[7]; however, it does not incorporate variational inference in its design. The aforementioned process parameterizes $p_{i}^{\prime}(z|y)$ as a multivariate Gaussian distribution and outlines the computation of $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$. We feed 8 vectors $\{z_{ij^{\prime}}\}^{14}_{j^{\prime}=7}$ into this distribution for calculating 8 logical probabilities $\{{p}_{i}^{\prime}(z_{ij^{\prime}}|\{z_{ij}\}^{6}_{j=1},\omega)\}_{j^{\prime}=7}^{14}$ in one perspective. In this paper, we calculate the final probability ${p}_{i}^{\prime}(z^{\prime}_{ij^{\prime}})$ by averaging the logical probabilities ${p}_{i}^{\prime}(z_{ij^{\prime}}|\{z_{ij}\}^{6}_{j=1},\omega)$ obtained from all perspectives. By utilizing the cross entropy loss function, we constrain the final probability ${p}_{i}^{\prime}(z^{\prime}_{i7})$ to approach 1 while ensuring other probabilities $\{{p}_{i}^{\prime}(z^{\prime}_{ij})\}_{j=8}^{14}$ are close to 0. Figure 6 illustrates in detail the forward process of the network $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ under a single perspective in $z^{\prime}_{ij}\in R^{n\times d}$. However, this approach did not yield satisfactory results.
@@
We speculate that the reason for the ineffectiveness of the aforementioned method lies in the unknown true form of the distribution $p_{i}^{\prime}(z|y)$, which may be a complex mixed Gaussian distribution. Therefore, it may not be reasonable to parameterize $p_{i}^{\prime}(z|y)$ as a multivariate Gaussian distribution and compute the probability $p_{i}^{\prime}(z_{ij^{\prime}}|y)$. Additionally, the formula for calculating the logical probability of a vector $z$ belonging to the parameterized distribution is provided as follows:
@@
Where $m$ represents the dimension index of the vector. $\mu$ represents the mean of the multivariate Gaussian distribution, $\sigma$ represents the standard deviation of the multivariate Gaussian distribution, and $z$ is the representation of some unknown data. It is evident that optimizing the network using the given formula (3) as the loss function presents challenges, as optimizing the above equation can be indirectly viewed as optimizing the Mean Squared Error (MSE) loss between vectors. Moreover, these vectors have not undergone normalization or squashing, making it difficult to optimize the network from the perspective of implementation.
To counter such ineffectiveness, in this paper, we train a deep network to directly fit probability $p_{i}^{\prime}(z_{ij^{\prime}}|y)$. Inevitably the network’s formulation has been transformed from $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ to $g_{\omega}(p_{i}^{\prime}(z_{ij^{\prime}}|y)|\{z_{ij}\}^{6}_{j=1},z_{ij^{\prime}})$, thereby simplifying the problem. Different from previous approaches $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ that fit the distribution $p_{i}^{\prime}(z|y)$, this paper employs a Transformer-Encoder to fit the probability $p_{i}^{\prime}(z_{ij^{\prime}}|y)$ directly, which enables the direct computation of the probability that $z_{ij^{\prime}}$ belongs to the distribution $p_{i}^{\prime}(z|y)$.
@@
For the implementation on the new form of $g_{\omega}(\cdot|\cdot)$, the function $g_{\omega}(\cdot|\cdot)$ maintains the egalitarian approach in processing each perspective $z_{ij}\in R^{d}$ within the outcomes $z^{\prime}_{ij}\in R^{n\times d}$ encoded by $f_{\theta}(z_{ij}|x_{ij})$. Within the confines of a singular perspective, we amalgamate $\{z_{ij}\}^{6}_{j=1}$ with one of the pending representations $\{z_{ij^{\prime}}\}^{14}_{j^{\prime}=7}$, resulting in a set of seven tokens. Following the embedding of positional encoding, these tokens are then systematically fed into a Transformer-Encoder, aimed at yielding 7 attention results. All 7 attention results are mapped to 7 probabilities by a MLP. The mean of the mapped outcomes is considered as the logical probability ${p}_{i}^{\prime}(z_{ij^{\prime}}|y)$ corresponding to the specified perspective. This constitutes the feedforward process of the $g_{\omega}(p_{i}^{\prime}(z_{ij^{\prime}}|y)|\{z_{ij}\}^{6}_{j=1},z_{ij^{\prime}})$. Each pending representation $z_{ij^{\prime}}$ in $\{z_{ij^{\prime}}\}^{14}_{j^{\prime}=7}$, receives a corresponding probability output $p_{i}^{\prime}(z_{ij^{\prime}}|y)$. Each logical probability stored in $\{{p}_{i}^{\prime}(z_{ij^{\prime}}|y)\}_{j^{\prime}=7}^{14}$ is obtained based on a single perspective. By averaging the logical probabilities across all perspectives, we derive the final probability values $\{p_{i}^{\prime}(z^{\prime}_{ij^{\prime}}|y)\}_{j^{\prime}=7}^{14}$. In this paper, we still optimize the model using the cross-entropy loss function. Figure 7 visually depicts the process of $g_{\omega}(p_{i}^{\prime}(z_{ij^{\prime}}|y)|\{z_{ij}\}^{6}_{j=1},z_{ij^{\prime}})$ handling a single perspective.
@@
In summary, this paper makes two compromises for PMoC. Firstly, we acknowledge that compared to $p_{i}^{\prime}(z)$, $q_{i}^{\prime}(z)$ is more likely to be a complex distribution. For instance, $p_{i}^{\prime}(z)$ may follow a Gaussian distribution, while $q_{i}^{\prime}(z)$ may adhere to a mixed Gaussian distribution. Therefore, estimating these two distributions and utilizing them to solve the Bongard-Logo problem is challenging to achieve excellent performance. Accordingly, we shift our focus to estimating the distribution $p_{i}^{\prime}(z)$ and utilize the probability of a representation belonging to this distribution to determine whether it is a primary or auxiliary group representation. Secondly, due to the challenges associated with optimizing the model from the perspective of estimating $p_{i}^{\prime}(z)$, this paper opts to directly compute the probability of a representation belonging to the distribution $p_{i}^{\prime}(z)$.
IV Methodology for abstract reasoning